# Good Noise Makes Good Edits: A Training-Free Diffusion-Based Video Editing with Image and Text Prompts

### v1

@@ Appendix B Additional quantitative results
CLIP-T $\uparrow$ DINO $\uparrow$ Temp $\uparrow$ PSNR $\uparrow$
FateZero fatezero 0.317 0.460 0.960 21.32
Video-p2p videop2p 0.300 0.393 0.961 21.75
Make-a-Pro make_a_pro 0.294 0.362 0.958 15.40
AnimateDiff animatediff + IP-Adapter ipadapter 0.302 0.361 0.958 24.83
ImEdit (Ours) 0.326 0.474 0.966 23.59
As shown in Table 3, ImEdit outperforms Video-P2P videop2p , Make-a-Pro make_a_pro , and FateZero fatezero across all metrics. Notably, in terms of DINO scores, our approach achieves higher reference similarity not only compared to text-based methods such as Video-P2 videop2p and FateZero fatezero but also surpasses image-conditioned methods like Make-a-Pro make_a_pro . While AnimateDiff + IP achieves a higher PSNR than ours, this difference stems from the smaller edited region in its outputs. ImEdit employs dilated dual masking during editing, which results in a broader generation area and a slightly lower PSNR.
@@ C.1 Image and Video Generation
Recent advancements in latent diffusion models (LDMs) have significantly improved the fidelity and controllability of image generation. Text-to-image (T2I) models such as Stable Diffusion stablediffusion , DALL-E 2 dalle2 , and Imagen imagen achieve high-resolution image synthesis aligned with complex text prompts. These models operate in the latent space for computational efficiency and support diverse conditioning mechanisms, including text, layout, and edge maps ipadapter ; controlnet . Their flexibility has made LDMs foundational components for extending generative modeling into the video domain.
Motivated by T2I success, text-to-video (T2V) models have emerged to generate coherent video sequences conditioned on textual descriptions. CogVideo cogvideo adapts T2I backbones with temporal attention, enabling efficient video generation with minimal overhead. ZeroScope zeroscope , built on ModelScope modelscope , improves visual fidelity and consistency, producing higher-resolution outputs while mitigating common artifacts in earlier methods. AnimateDiff animatediff integrates temporal modules into Stable Diffusion stablediffusion , facilitating the T2V model with improved coherence between frames.
Alongside architectural advances, controllability and identity preservation have gained importance. Several methods leverage a reference frame along with a prompt to guide spatial appearance and layout across time stiv ; ti2v ; controlvideo . In addition, others introduce subject-driven personalization via instance finetuning videoalchemist ; dreamvideo ; tune_a_video ; customcrafter , enabling video generation with identity consistency across diverse prompts. These trends illustrate a broader shift toward multimodal conditioning across text, image, and motion to allow more expressive and controllable video synthesis.
@@ C.2 Image and Video Editing
Following the advances in image and video generation models, diffusion models have also made significant progress in image and video editing, allowing intuitive and precise content manipulation. In the image domain, Prompt-to-Prompt prompt-to-prompt introduces cross-attention manipulation within pretrained diffusion models, enabling semantic edits without finetuning or masks. InstructPix2Pix instructpix2pix leverages pairs of instructions and images to support editing guided by natural language. EditAnything editanything generalizes this paradigm by integrating multi-modal inputs such as text, masks, and user clicks.
Video editing extends these capabilities to the temporal dimension. Tune-A-Video tune_a_video employs one-shot tuning of pretrained image diffusion models for video editing tasks, allowing users to modify video content and style with minimal data and strong temporal consistency. Video-P2P videop2p extends cross-attention control techniques from image to video, maintaining both spatial and temporal coherence. Text2Video-Zero text2video removes the need for paired text-video data, instead leveraging pretrained text-image models for flexible and data-efficient video editing.
To overcome the limitations of editing based solely on text, recent studies propose image-conditioned video editing methods that use a reference image to guide appearance and identity genvideo ; make_a_pro ; cutandpaste . Make-A-Protagonist make_a_pro introduces a unified framework for protagonist replacement, background editing, and style transfer using mask-guided denoising. GenVideo genvideo proposes shape-aware masking and latent noise correction to ensure temporal consistency and subject fidelity. VideoSwap videoswap adds semantic point correspondences to provide interactive and precise control over the subject’s motion and appearance. Despite these advancements, existing methods still require a finetuning stage with the source video, which limits their overall effectiveness. In contrast, our approach eliminates this need entirely, enabling high-quality, zero-shot video editing with strong temporal and identity coherence.
@@ Appendix D Limitation
Although our proposed method exhibits notable improvements over existing baselines, as evidenced by extensive experiments, it still faces certain limitations. When there is a substantial disparity in the innate physical characteristics between the source and target objects, our method struggles to perform object replacement. Figure 11 shows that our method generates wings for the red panda which is an anatomically implausible feature to transfer the parrot’s flying motion. Conversely, in attempting to preserve the red panda’s natural characteristics, the method fails to convey the intended motion.
Figure 11: Example of failure case. A large semantic gap between the object in the source video and the reference image poses challenges for our method. We show results from two versions of our method, obtained by varying the sampling start point $\rho$ and the dilation value. In this case, editing a bird into a red panda results in missing or incorrect transfer of the wings and arms.
Figure 12: Additional results on the proposed dilated dual masking. (a) shows the output of the first stage using the source video $\mathcal{V}_{s}$ and source object mask, while (b)-(d) present outputs of the second stage from the intermediate video $\mathcal{V}_{r}$ using different masking strategies. Masks are overlaid on the corresponding video. The proposed dilated dual masking (d) achieves the most accurate edits, underscoring its effectiveness for the object replacement.
Appendix E Additional ablation results on dilated dual masking
The proposed dilated dual masking enables the seamless replacement of the source object with the target object from the reference image across all frames. Fig. 12 presents the complete results of Fig. 7, confirming that the patterns observed remain consistent across different frames.
In Fig. 12(a), when only the source mask is used, the target object is unnaturally constrained to the shape of the source mask, resulting in distorted generations. In Fig. 12(b), using only the rough mask fails to fully cover the source object, leaving undesired remnants of it in the edited video. In Fig. 12(c), employing a dual mask obtained by taking the union of the source and target masks alleviates artifacts from the source object and shape mismatches of the target object, but the mask boundaries still lack smoothness. In contrast, our proposed dilated dual mask in Fig. 12(d) effectively mitigates these issues, enabling natural removal of the source object and coherent synthesis of the target object throughout the video.
Appendix F Additional experimental settings
@@
Appendix F.1 provides additional implementation details, including hyperparameter settings. Appendix F.2 describes the evaluation questions used in our user study.
F.1 Additional implementation details
Our method involves four key hyperparameters that control different aspects of the video editing process, which include $\rho,\gamma,\text{dilation kernel size $k$, and guidance scale}$:
  • $\rho\in[10,30]$: Control the starting timestep for the sampling process in $\rho$-start sampling, determining the trade-off between motion preservation from the source video and appearance transformation toward the reference image.
  • $\gamma\in[0,1]$: Control the weighted combination of standard zero-vector and zero-image embeddings in zero image guidance, enabling fine-grained adjustment of color fidelity in the generated object.
  • dilation kernel size $k$ $\in[0,10]$: Specifies the size of the dilation kernel used in the proposed dilated dual masking. This affects the spatial smoothness and coverage of the edited region.
  • guidance scale $\in[6,10]$: Determine the strength of classifier-free guidance during denoising. A higher guidance scale enhances fidelity to the conditioning signal.
F.2 User study details
@@
In the user study described in Section  3.2.1, participants are asked to evaluate five key aspects of the generated videos using the following ranking-based questions:
@@
  • Image Alignment: Rank the videos from 1st to 4th based on how well the appearance of the "object in the generated video" matches the "object in the reference image (excluding the background)".
@@
  • Text Alignment: Rank the videos from 1st to 4th based on how well the "generated video content" aligns with the "provided text prompt".
@@
  • Motion Alignment: Rank the videos from 1st to 4th based on how closely the "motion of the object in the generated video" matches the "motion of the object in the source video".
  • Temporal Consistency: Rank the videos from 1st to 4th based on the "visual consistency of the object and background throughout the edited video".
  • Overall Quality: Taking all the above aspects into consideration, rank the videos from 1st to 4th based on the overall realism and naturalness.
Appendix G Algorithm
The overall video editing procedure of ImEdit is summarized in Algorithm 1. Our training-free method consists of two stages: a rough editing stage that generates an initial object transformation guided by the source mask, and a refinement stage that further improves visual fidelity and temporal coherence using an updated mask derived from the intermediate result.
Algorithm 1 Training-free video editing procedure of ImEdit
@@ Appendix G Algorithm
4:${\bm{c}}_{\text{img}}\leftarrow\text{CLIP}_{\text{img}}(\mathcal{R}),{\bm{c}}_{\text{txt}}\leftarrow\text{CLIP}_{\text{txt}}(\mathcal{T})$
5:$\bar{{\bm{c}}}_{\text{img}}=\gamma\cdot\text{CLIP}_{\text{img}}({\mathbf{I}}_{0})+(1-\gamma)\cdot\mathbf{0}$, $\bar{{\bm{c}}}_{\text{txt}}=\gamma\cdot\text{CLIP}_{\text{txt}}({\mathcal{\bar{T}}}),\text{ where }\gamma\in[0,1]$
@@
8:${\bm{M}}_{s}\leftarrow\text{Segment}(\mathcal{V}_{s})$
9:${\bm{M}}^{\prime}_{s}\leftarrow\text{Dilate}({\bm{M}}_{s})$
@@
11:${\bm{z}}_{0}\leftarrow\mathcal{E}(\mathcal{V}_{s})$
12:$\{{\bm{z}}_{t}\}_{t=\tau_{1}}^{\tau_{\nu}}\leftarrow\text{DDIM-inv}({\bm{z}}_{0})$
@@
15:${\bm{z}}^{\prime}_{\tau_{\rho}}\leftarrow\sqrt{\alpha_{\tau_{\rho}}}\cdot{\bm{z}}_{0}+\sqrt{1-\alpha_{\tau_{\rho}}}\cdot\eta$, where $\eta\sim\mathcal{N}(0,{\bm{I}})$
16:for $t=\rho$ to $1$ do
@@
22:$\mathcal{V}_{r}\leftarrow\mathcal{D}({\bm{z}}^{\prime}_{\tau_{0}})$
23:$\mathcal{V}_{r}\leftarrow\mathcal{V}_{r}\odot{\bm{M}}^{\prime}_{s}+\mathcal{V}_{s}\odot(1-{\bm{M}}^{\prime}_{s})$
@@
26:${\bm{M}}_{r}\leftarrow\text{Segment}(\mathcal{V}_{r})$
27:${\bm{M}}^{\prime}_{r}\leftarrow\text{Dilate}({\bm{M}}_{r})$
28:${\bm{M}}_{e}\leftarrow{\bm{M}}^{\prime}_{s}\cup{\bm{M}}^{\prime}_{r}$
@@
30:${\bm{z}}_{0}\leftarrow\mathcal{E}(\mathcal{V}_{r})$
31:$\{{\bm{z}}_{t}\}_{t=\tau_{1}}^{\tau_{\nu}}\leftarrow\text{DDIM-inv}({\bm{z}}_{0})$
@@
34:${\bm{z}}^{\prime}_{\tau_{\rho}}\leftarrow\sqrt{\alpha_{\tau_{\rho}}}\cdot{\bm{z}}_{0}+\sqrt{1-\alpha_{\tau_{\rho}}}\cdot\eta$, where $\eta\sim\mathcal{N}(0,{\bm{I}})$
35:for $t=\rho$ to $1$ do
@@
41:$\mathcal{V}_{e}\leftarrow\mathcal{D}({\bm{z}}^{\prime}_{\tau_{0}})$
Appendix H Additional qualitative results
Notably, AnimateDiff+IP-Adapter frequently exhibits noticeable distortions in both shape and color, leading to unnatural visual artifacts. To address this, our method constructs a well-formed noise maps by selecting a less-degraded latent via $\rho$-start sampling and combining foreground and background latents through dilated dual masking. This structured noise formulation enables accurate object replacement while preserving the motion dynamics of the source video.
Figure 13: Complete video editing comparisons with baseline methods. FateZero fatezero and Make-a-Pro make_a_pro use text prompts generated by BLIP-2 blip2 , following their respective proposed strategies. Video-P2P videop2p and FateZero fatezero use only text prompts, while Make-a-Pro make_a_pro , AnimateDiff animatediff +IP-Adapter ipadapter , and our proposed method use both image and text prompts.
Figure 14: Complete video editing comparisons with baseline methods. FateZero fatezero and Make-a-Pro make_a_pro use text prompts generated by BLIP-2 blip2 , following their respective proposed strategies. Video-P2P videop2p and FateZero fatezero use only text prompts, while Make-a-Pro make_a_pro , AnimateDiff animatediff +IP-Adapter ipadapter , and our proposed method use both image and text prompts.
Figure 15: Complete video editing comparisons with baseline methods. FateZero fatezero and Make-a-Pro make_a_pro use text prompts generated by BLIP-2 blip2 following their respective proposed strategies. Video-P2P videop2p and FateZero fatezero use only text prompts, while Make-a-Pro make_a_pro , AnimateDiff animatediff +IP-Adapter ipadapter , and our proposed method use both image and text prompts.

### v2

@@ Appendix B Full qualitative comparison and ablations
DINO $\uparrow$ CLIP-T $\uparrow$ Temp $\uparrow$ PSNR $\uparrow$
AnimateDiff [animatediff] 0.412 0.308 0.980 13.9
+ Ours 0.498 0.321 0.972 24.7
TokenFlow [tokenflow] 0.480 0.325 0.975 19.6
+ Ours 0.523 0.320 0.974 23.9
B.1 Qualitative comparison with baselines
As shown in Fig. 10, since SDEdit-Video [sdedit] and AnimateDiff [animatediff]+IP-Adapter [ipadapter] are adapted or composed from methods not originally designed for video editing, their qualitative results exhibit reduced visual fidelity compared to other baseline methods. Specifically, SDEdit-Video is extended from its original image-based formulation to the video domain, while AnimateDiff+IP-Adapter combines AnimateDiff with IP-Adapter in a naive manner. As a result, SDEdit-Video preserves only the coarse spatial layout of the source video and often introduces excessive noise, leading to degraded visual quality. In the case of AnimateDiff+IP-Adapter, reference images are directly pasted into the edited regions without sufficient consideration of semantic structure, resulting in visually unnatural artifacts.
B.2 Qualitative ablation studies
Dilated dual masking. Dilated dual masking enables the seamless replacement of the source object with the target object in the reference image across frames. Fig. 11 presents the complete results of Fig. 7, confirming that the patterns observed remain consistent across different frames.
In Fig. 11(a), when only the source mask is used, the target object is unnaturally constrained to the shape of the source mask, resulting in distortion. In Fig. 11(b), using only the rough mask fails to fully cover the source object, leaving undesired remnants of it in the edited video. In Fig. 11(c), employing a dual mask obtained by taking the union of the source and target masks, alleviates artifacts from the source object and shape mismatches of the target object, but the mask boundaries still lack smoothness. In contrast, our dilated dual mask in Fig. 11(d) effectively mitigates these, enabling clean removal of the source object and coherent synthesis of the target object throughout the video.
$\rho$-start sampling. Our $\rho$-start sampling initiates DDIM sampling and inversion from intermediate time steps, thereby balancing the appearance of the reference image with the spatial layout of the source video. Fig. 12 supplements Fig. 8 by additionally presenting the final edited video results. Specifically, while Fig. 8 only illustrates the first stage edits under varying values of $\rho$, the rightmost column in Fig. 12 provides the corresponding final outputs after the second stage. The second stage further refines facial details, producing results that more closely resemble the reference image.
Zero image guidance. Zero image guidance is introduced specifically to handle the occasional color desaturation observed in the edited videos. By adjusting the negative prompt to steer the generation away from the zero image direction, this mechanism effectively enhances color fidelity. Figure 13 provides an extension of the results in Figure 9 by presenting edits across a wider range of $\gamma$ values. As $\gamma$ approaches 1, the contribution of zero image guidance within the negative image guidance increases, leading to noticeably improved color saturation in the edited videos.
Figure 12: Full qualitative results of the ablation study on the $\rho$-start sampling. We present the complete results corresponding to Fig. 8 in the main manuscript by adding final edited results. As the value of $\rho$ increases from 10 to 40 in Stage 1, the parrot in the source video gradually transforms to more closely resemble the owl in the reference image. After processing through Stage 2, the final edits further align with the reference appearance and exhibit improved visual quality.
Figure 13: Full qualitative results of the ablation study on the zero image guidance scale $\gamma$ . We include the complete results corresponding to Fig. 9 in the main paper. As the $\gamma$ value is adjusted, the color fidelity of the edited video is calibrated to better match the reference image.
Figure 14: Example of failure case. A large semantic gap between the object in the source video and the reference image poses challenges for our method. We show results from two versions of our method, obtained by varying the sampling start point $\rho$ and the dilation kernel size. In this case, editing a bird into a red panda results in missing or incorrect transfer of the wings and arms.
@@ C.1 Image and video generation
Recent advancements in latent diffusion models (LDMs) have significantly improved the fidelity and controllability of image generation. Text-to-image (T2I) models such as Stable Diffusion [stablediffusion], DALL-E 2 [dalle2], and Imagen [imagen] achieve high-resolution image synthesis aligned with complex text prompts. These models operate in the latent space for computational efficiency and support diverse conditioning mechanisms, including text, layout, and edge maps [ipadapter, controlnet]. Their flexibility has made LDMs foundational components for extending generative modeling into the video domain.
Motivated by T2I success, text-to-video (T2V) models have emerged to generate coherent video sequences conditioned on textual descriptions. CogVideo [cogvideo] adapts T2I backbones with temporal attention, enabling efficient video generation with minimal overhead. ZeroScope [zeroscope], built on ModelScope [modelscope], improves visual fidelity and consistency, producing higher-resolution outputs while mitigating common artifacts in earlier methods. AnimateDiff [animatediff] integrates temporal modules into Stable Diffusion [stablediffusion], facilitating the T2V model with improved coherence between frames.
Alongside architectural advances, controllability and identity preservation have gained importance. Several methods leverage a reference frame along with a prompt to guide spatial appearance and layout across time [stiv, ti2v, controlvideo]. In addition, others introduce subject-driven personalization via instance finetuning [videoalchemist, dreamvideo, tune_a_video, customcrafter], enabling video generation with identity consistency across diverse prompts. These trends illustrate a broader shift toward multimodal conditioning across text, image, and motion to allow more expressive and controllable video synthesis.
@@ C.2 Image and video editing
Following the advances in image and video generation models, diffusion models have also made significant progress in image and video editing, allowing precise content manipulation. In the image domain, Prompt-to-Prompt [prompt-to-prompt] introduces cross-attention manipulation within pretrained diffusion models, enabling semantic edits without finetuning or masks. InstructPix2Pix [instructpix2pix] leverages pairs of instructions and images to support editing guided by natural language. EditAnything [editanything] generalizes this paradigm by integrating multi-modal inputs such as text, masks, and user clicks.
Video editing extends these capabilities to the temporal dimension. Tune-A-Video [tune_a_video] employs one-shot tuning of pretrained image diffusion models for video editing tasks, allowing users to modify video content and style with minimal data and strong temporal consistency. Video-P2P [videop2p] extends cross-attention control techniques from image to video, maintaining both spatial and temporal coherence. Text2Video-Zero [text2video] removes the need for paired text-video data, instead leveraging pretrained text-image models for flexible and data-efficient video editing.
To overcome the limitations of editing based solely on text, recent studies propose image-conditioned video editing methods that use a reference image to guide appearance and identity [genvideo, make_a_pro, cutandpaste]. Make-A-Protagonist [make_a_pro] introduces a unified framework for protagonist replacement, background editing, and style transfer using mask-guided denoising. GenVideo [genvideo] proposes shape-aware masking and latent noise correction to ensure temporal consistency and subject fidelity. VideoSwap [videoswap] adds semantic point correspondences to provide interactive and precise control over the subject’s motion and appearance. Despite these advancements, existing methods still require finetuning on each source video and rely on user-defined semantic keypoints, limiting practicality and scalability. In contrast, our method directly utilizes the image prompt without any finetuning, while effectively preserving temporal consistency and semantic accuracy.
@@ Appendix D Limitation
Although our method exhibits notable improvements over existing baselines, as evidenced by extensive experiments, it still faces certain limitations. When there is a substantial disparity in the innate physical characteristics between the source and target objects, our method struggles to perform object replacement. Figure 14 shows that our method generates wings for the red panda which is an anatomically implausible feature to transfer the parrot’s flying motion. Conversely, in attempting to preserve the red panda’s natural characteristics, it fails to convey the intended motion.
Appendix E Additional experimental settings
@@
Appendix E.1 provides further implementation details, including hyperparameter settings. Appendix E.2 outlines the reproducibility settings for all baseline models used in our experiments. Appendix E.3 presents the evaluation questions used in our user study.
E.1 Additional implementation details
Our method introduces four key hyperparameters: $\rho$, $\gamma$, dilation kernel size, and guidance scale. Each parameter has a well-defined and narrow range that yields consistent performance across diverse editing tasks. Consequently, hyperparameter tuning in the proposed method requires minimal manual adjustment and can be efficiently performed without extensive search or optimization.
  • $\rho\in[10,30]$: Controls the starting timestep for the denoising trajectory. In practice, one of the discrete values (10, 20, or 30) generally performs well. Users may optionally refine the result by exploring intermediate values.
  • $\gamma\in[0,1]$: Controls the weighted combination of standard zero-vector and zero-image embeddings in zero image guidance. The default value of 0.5 has been empirically found to perform robustly across a wide variety of scenes, object types, and motion dynamics.
  • dilation kernel size $k$ $\in[0,10]$: Defines the kernel size used in the dilated dual masking. It affects the spatial smoothness and coverage of the edited region. The default value of 3 works well in most cases, while slightly larger values (e.g., 5 or 7) can help when the source and target objects differ significantly in shape or size.
  • guidance scale $\in[6,10]$: Determines the strength of classifier-free guidance during denoising. The default setting of 6 typically yields semantically faithful results.
Depending on the resolution of the source video, we use one of the following sizes: 320×576, 432×768, 480×480, 515×512, 600×600, 640×640, or 700×700. As an exception, the results in Fig. LABEL:fig:teaser are rendered at 1024×576 to better showcase the our method’s performance. The number of frames is set to 8 or 16, selected according to the source video length.
E.2 Reproducibility of baseline models
All baseline models are run using the official implementations released in their GitHub repositories. For a fair comparison, we evaluate all baselines under the same experimental settings as our method unless specified otherwise (refer to Appendix E.1). The same text prompts are applied across all methods, except for the method that follows its own text prompting strategy. We also use the same spatial resolution for VINO and all baseline models, except for methods that inherently require square inputs.
  • SDEdit-Video [sdedit] [^1] https://github.com/huggingface/diffusers/blob/main/src/diffusers/pipelines/animatediff/pipeline_animatediff_video2video.py We adopt the video-extended implementation provided by HuggingFace, as the original SDEdit is originally image-based.
  • FateZero [fatezero] [^2] https://github.com/ChenyangQiQi/FateZero As FateZero [fatezero] supports only square outputs, we paste its generated video into the corresponding region of the source video for comparison. Both FateZero [fatezero] and Make-A-Pro [make_a_pro] employ model-based prompting procedures to automatically generate text prompts. While FateZero originally uses an earlier BLIP variant for this step, the overall prompting procedure is fundamentally the same as that of Make-A-Protagonist, which uses BLIP-2. To maintain consistency, we adopt the more recent BLIP-2–based prompting strategy for FateZero as well and use identical text prompts for both methods.
  • Video-P2P [videop2p] [^3] https://github.com/dvlab-research/Video-P2P Similar to FateZero, since Video-P2P [videop2p] produces square outputs, we paste its generated video into the corresponding region of the source video.
  • TokenFlow [tokenflow] [^4] https://github.com/omerbt/TokenFlow TokenFlow is evaluated under the same experimental settings as our method without any additional modifications. To implement TokenFlow+Ours, we extend TokenFlow by adding IP-Adapter [ipadapter] to support image conditioning, and incorporate our proposed approaches, including zero-image guidance, $\rho$-start sampling, and dilated dual masking.
  • AniamteDiff [animatediff] [^5] https://github.com/guoyww/AnimateDiff+IP-Adapter [ipadapter] [^6] https://github.com/tencent-ailab/IP-Adapter For both AnimateDiff and IP-Adapter, we use their official GitHub implementations, and fuse the two following the integration scheme proposed in the IP-Adapter paper.
  • Make-A-Protagonist [make_a_pro] [^7] https://github.com/HeliosZhao/Make-A-Protagonist Following the original prompt generation strategy, we extract protagonist cues using BLIP-2 [blip2] and generate prompts via visual question answering as provided in its official GitHub repository.
E.3 User study details
@@
In the user study described in Section 3.1, participants were asked to rank the seven generated videos from 1st (best) to 7th (worst) for each of the following five evaluation criteria:
@@
  • Image Alignment: How well the appearance of the "object in the generated video" matches the "object in the reference image (excluding the background)".
@@
  • Text Alignment: How well "the content of the generated video" aligns with the "provided text prompt".
@@
  • Motion Alignment: How closely the "motion of the object in the generated video" matches the "motion of the object in the source video".
  • Temporal Consistency: The degree to which the object and background remain consistent across time.
  • Overall Quality: The overall realism and naturalness of the edited video, considering all aspects above.
Appendix F Algorithm
The overall video editing pipeline of VINO is outlined in Algorithm 1. Our training-free method consists of two stages: a rough editing stage that generates an initial object transformation guided by the source mask, and a refinement stage that further improves visual fidelity and temporal coherence using an updated mask derived from the intermediate result.
Algorithm 1 Training-free video editing procedure of VINO
@@ Appendix F Algorithm
[…]
@@
[…]
@@
[…]
@@
[…]
16:for $t=\rho$ to 11 do
@@
[…]
@@
[…]
@@
[…]
@@
[…]
35:for $t=\rho$ to 11 do
@@
[…]
Appendix G Additional qualitative results
Additional qualitative comparisons with baseline methods are provided in Fig. 15, Fig. 16, Fig. 17, Fig. 18, and Fig. 19, which further demonstrate the superior visual fidelity and temporal consistency achieved by our approach. Notably, AnimateDiff+IP-Adapter frequently exhibits noticeable distortions in both shape and color, leading to unnatural visual artifacts. To address this, our method constructs a well-formed noise maps by selecting a less-degraded latent via $\rho$-start sampling and combining foreground and background latents through dilated dual masking. This structured noise formulation enables accurate object replacement while preserving the motion dynamics of the source video.
Figure 15: Complete video editing comparisons with baseline methods. FateZero [fatezero] and Make-a-Pro [make_a_pro] use text prompts generated by BLIP-2 [blip2], following their respective proposed strategies. TokenFlow [tokenflow], Video-P2P [videop2p], FateZero [fatezero], and SDEdit-Video [sdedit] use only text prompts, while Make-a-Pro [make_a_pro], AnimateDiff [animatediff]+IP-Adapter [ipadapter], and our proposed method use both image and text prompts.
Figure 16: Complete video editing comparisons with baseline methods. FateZero [fatezero] and Make-a-Pro [make_a_pro] use text prompts generated by BLIP-2 [blip2] following their respective proposed strategies. TokenFlow [tokenflow], Video-P2P [videop2p], FateZero [fatezero], and SDEdit-Video [sdedit] use only text prompts, while Make-a-Pro [make_a_pro], AnimateDiff [animatediff]+IP-Adapter [ipadapter], and our proposed method use both image and text prompts.
Figure 17: Complete video editing comparisons with baseline methods. FateZero [fatezero] and Make-a-Pro [make_a_pro] use text prompts generated by BLIP-2 [blip2] following their respective proposed strategies. TokenFlow [tokenflow], Video-P2P [videop2p], FateZero [fatezero], and SDEdit-Video [sdedit] use only text prompts, while Make-a-Pro [make_a_pro], AnimateDiff [animatediff]+IP-Adapter [ipadapter], and our proposed method use both image and text prompts.
Figure 18: Complete video editing comparisons with baseline methods. FateZero [fatezero] and Make-a-Pro [make_a_pro] use text prompts generated by BLIP-2 [blip2] following their respective proposed strategies. TokenFlow [tokenflow], Video-P2P [videop2p], FateZero [fatezero], and SDEdit-Video [sdedit] use only text prompts, while Make-a-Pro [make_a_pro], AnimateDiff [animatediff]+IP-Adapter [ipadapter], and our proposed method use both image and text prompts.
Figure 19: Complete video editing comparisons with baseline methods. FateZero [fatezero] and Make-a-Pro [make_a_pro] use text prompts generated by BLIP-2 [blip2] following their respective proposed strategies. TokenFlow [tokenflow], Video-P2P [videop2p], FateZero [fatezero], and SDEdit-Video [sdedit] use only text prompts, while Make-a-Pro [make_a_pro], AnimateDiff [animatediff]+IP-Adapter [ipadapter], and our proposed method use both image and text prompts.